OpenStudy (anonymous):

On Tuesday 84 of the 126 students usually attend school were unable to attend due to heavy snowfall. What is the GCF of these two numbers

OpenStudy (lalaly):

42

OpenStudy (anonymous):

How did you figure it.

OpenStudy (anonymous):

Break each number down into their prime factors. 84 = 2 * 2 * 3 * 7. 126 = 2 * 3 * 3 * 7. Multiply all the common factors together. 2 is a common factor, 3 is a common factor, and 7 is a common factor. 2 * 3 * 7 = 42 is the GCF.

OpenStudy (anonymous):

thank you